Circular No. 6222

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1735–27 AND GRO J1750–27 T. Koh, D. Chakrabarty, T. A. Prince, and B. Vaughan, Califor-

nia Institute of Technology; and S. N. Zhang, M. Scott, M. H. Finger, and R. B. Wilson, Marshall Space Flight Center, report for the Compton GRO/BATSE team: “Further observations of the recently-discovered, 4.45s transient hard x-ray pulsar in the Galactic center region (GRO J1735–27; see IAUC 6207) have yielded a refined position, based only on pulsed flux, which indicates that the pulsed and unpulsed emission reported on IAUC 6207 come from different sources. The estimated 90-percent-confidence error box for the pulsed source is approximately a quadrilateral of size 0.7 deg^{2}, with vertices α = 17^{h}47^{m}, δ = −27^{o}.8; 17^{h}50^{m}, −28^{o}.0; 17^{h}52^{m}, −26^{o}.9; 17^{h}49^{m}, −26^{o}.7 (equinox 2000.0). In accordance with the new position, we have designated the pulsed source as GRO J1750–27, the previous designation (GRO J1735–27) referring to the unpulsed source. The pulsed emission from GRO J1750–27, first detected on July 29 UT, continues to be visible as of Aug. 29; its barycentric pulse period on Aug. 29.5 was 4.45311 ± 0.00003 s, with a period derivative of ∼ −0.11 s/yr. Optical and soft-x-ray observations of the error box are encouraged to constrain further the source position and to identify the accreting companion. The unpulsed source (GRO J1735–27) is not observed in data combined during Aug. 11–21, while pulses from GRO J1750–27 have been detected at a nearly constant level through Aug. 29. The position reported on IAUC 6207 remains valid for the unpulsed source GRO J1735–27.”

SUPERNOVA 1995Y IN NGC 410 M. Cavagna, Sormano, Italy, reports the following precise position for

SN 1995Y (cf. IAUC 6221) from a CCD image obtained on Sept. 1.945 UT, when the supernova was at V = 18: α = 1^{h}10^{m}59^{s}.60, δ = +33^{o}08′40′′.7 (equinox 2000.0; GSC reference stars); this corresponds to a point 8′′.4 east and 26′′.1 south of the galaxy’s center. M. Tombelli, Montelupo, Italy, reports the following position end figures from a CCD image obtained on Sept. 3.063: 59^{s}.70, 42′′.8 (also GSC reference stars).

SUPERNOVA 1995Z IN UGC 937 Cavagna also sends the following precise position of SN 1995Z from a

CCD image obtained on Sept. 1.954 UT: α = 1^{h}23^{m}38^{s}.35, δ = +32^{o}37′53′′.1 (equinox 2000.0; magnitude V = 19).

1995 September 6 Daniel W. E. Green